OpenStudy (anonymous):

Show how to solve: sin x = 1/2... The answer is this: pi/6, 5pi/6

OpenStudy (anonymous):

This involves the unit circle. You would have to know the unit circle to be able to say that the answers are pi/6 and 6pi/6. The most I can do is say look at the unit circle. Sine is equivalent to the y value. So when y = 1/2, what is the angle? 30 degrees and 150 degrees (which, when translated to radians, become the ones you got)

OpenStudy (anonymous):

1/2 = sin30 degree = sin pi/6

OpenStudy (anonymous):

oh thats why.. im confused though because of the x indicated in the question, so sin is y, the cos is x? and tan is?

OpenStudy (anonymous):

-> sin x = sin π/6 => x = π/6 From the unit circle: sin π/6 = sin ( π - π/6) => = sin 5π/6 Thus x = π/6, 5π/5

OpenStudy (anonymous):

@cdogg007 , tangent is y/x